# Towards Practical Benchmarking of Data Cleaning Techniques: On Generating Authentic Errors via Large Language Models

### v1

@@ 7. Related Work
Ensuring data quality in relational databases and tabular data processing remains a significant challenge, as errors like missing values, pattern violations, and rule violations can disrupt downstream analytical and machine learning tasks. Traditional rule-based approaches, such as functional dependencies (FDs), uniqueness constraints (UCs), and pattern constraints, have been widely used (Rekatsinas et al., 2017). However, these methods require extensive manual specification, limiting their adaptability to large-scale, high-dimensional data. Recent advancements in deep learning and large language models (LLMs) have introduced more flexible solutions for data quality management. Transformer-based models have shown strong generalization across various tasks, including data imputation, error detection, and error correction (Li et al., 2024a; Deng et al., 2022). AI-driven methods further enhance adaptability by learning complex data patterns and reducing manual intervention (Zhu et al., 2024).
@@ 7.1. Data Imputation
Data imputation is essential for maintaining data completeness. Traditional methods include statistical techniques (e.g., mean, median, mode (Wei et al., 2018)), matrix factorization (SVD (Troyanskaya et al., 2001), PCA (Oba et al., 2003)), regression-based approaches (LOESS (Cleveland and Loader, 1996), IIM (Zhang et al., 2019)), and nearest-neighbor imputation (kNN (Regression, 1992), MIBOS (Wu et al., 2012)). Constraint-based methods leverage data integrity constraints (MLEM2 (Grzymala-Busse et al., 2005), DERAND (Song et al., 2015b, 2018)), while multiple imputation techniques (MICE (Van Buuren and Groothuis-Oudshoorn, 2011), SICE (Khan and Hoque, 2020)) generate multiple candidate values to improve robustness. However, these approaches often rely on strong assumptions, limiting their adaptability to complex data structures.
Machine learning (ML) and deep learning (DL) have improved imputation by capturing intricate data dependencies. Neural network-based models like RNNI (Rumelhart et al., 1986) and MLPI (García-Laencina et al., 2010) enhance generalization, while autoencoders (MIDAE (Gondara and Wang, 2018), SDAE (Costa et al., 2018)), VAEs (HI-VAE (Nazabal et al., 2020)), and graph neural networks (GRAPE (You et al., 2020), IGRM (Zhong et al., 2023)) provide structural awareness. GAN-based approaches (GAIN (Yoon et al., 2018), WGAN (Arjovsky et al., 2017)) further refine imputation by generating plausible missing values. More recently, large language models (LLMs) like GPT-3 (Narayan et al., 2022) and Table-GPT (Li et al., 2024a) leverage zero-shot and few-shot learning for imputation, offering strong generalization but facing limitations in handling numerical data. Despite these advancements, DL-based methods remain computationally expensive and less interpretable than traditional approaches.
@@ 7.2. Error Detection
Error detection is essential for identifying anomalies in relational data. Traditional methods include statistical techniques (e.g., PCA (Shyu et al., 2003), dBoost (Pit-Claudel et al., 2016)) and constraint-based approaches enforcing integrity constraints (e.g., FDs (Mandros et al., 2017), CFDs (Fan et al., 2008), DCs (Chu et al., 2013a)). Additionally, clustering-based methods (e.g., DBSCAN (Ester et al., 1996), OPTICS (Ankerst et al., 1999)) and outlier detection techniques (e.g., LOF (Breunig et al., 2000), iForest (Liu et al., 2008)) have also been widely used to capture anomalies.
Beyond traditional approaches, machine learning (ML)-based approaches enhance detection by integrating multiple detection signals and classifier models (e.g., Raha (Mahdavi et al., 2019), HoloDetect (Heidari et al., 2019)). Unsupervised (e.g., Uni-Detect (Wang and He, 2019)) and active learning techniques (e.g., ED2 (Neutatz et al., 2019)) further improve adaptability by reducing reliance on labeled data. Deep learning (DL)-based methods advance error detection by employing autoencoders for anomaly detection (e.g., RandNet (Chen et al., 2017), REPEN (Pang et al., 2019)) and utilizing semi-supervised learning to model normal and erroneous data distributions (e.g., DeepSAD (Ruff et al., 2019)). Recently, large language models (LLMs) have been fine-tuned for tabular data anomaly detection (e.g., ADTDL (Li et al., 2024b)), showing potential despite computational challenges.
@@ 7.3. Error Repairing
Error repairing aims to restore data integrity by correcting erroneous values. Traditional methods include statistical reasoning (e.g., ERACER (Mayfield et al., 2010), SCARE (Yakout et al., 2013)) and constraint-based approaches enforcing FDs and CFDs (e.g., NADEEF (Ebaid et al., 2013), Holistic (Chu et al., 2013b)). Clustering-based strategies leverage data similarity for correction (e.g., DORC (Song et al., 2015a), DISC (Song et al., 2021)).
Recent ML-based methods integrate probabilistic reasoning and external data sources (e.g., HoloClean (Rekatsinas et al., 2017), BoostClean (Krishnan et al., 2017)) to refine repairs. DL-based approaches enhance repair by learning data dependencies (e.g., DeepClean (Zhang et al., 2018)) and leveraging large-scale contextual embeddings (e.g., LLMClean (Biester et al., 2024)). While these methods improve repair accuracy, they remain computationally expensive, requiring further optimization.

### v2

@@ 7. Related Work
Ensuring data quality in relational databases and tabular data processing remains a significant challenge, as errors like missing values, pattern violations, and rule violations can disrupt downstream analytical and machine learning tasks. Traditional rule-based approaches, such as functional dependencies (FDs), uniqueness constraints (UCs), and pattern constraints, have been widely used (rekatsinas2017holoclean). However, these methods require extensive manual specification, limiting their adaptability to large-scale, high-dimensional data. Recent advancements in deep learning and large language models (LLMs) have introduced more flexible solutions for data quality management. Transformer-based models have shown strong generalization across various tasks, including data imputation, error detection, and error correction (li2024table; deng2022turl). AI-driven methods further enhance adaptability by learning complex data patterns and reducing manual intervention (zhu2024relational).
@@ 7.1. Data Imputation
Data imputation is essential for maintaining data completeness. Traditional methods include statistical techniques (e.g., mean, median, mode (wei2018missing)), matrix factorization (SVD (troyanskaya2001missing), PCA (oba2003bayesian)), regression-based approaches (LOESS (cleveland1996smoothing), IIM (zhang2019learning)), and nearest-neighbor imputation (kNN (regression1992introduction), MIBOS (wu2012missing)). Constraint-based methods leverage data integrity constraints (MLEM2 (grzymala2005handling), DERAND (song2015enriching; song2018enriching)), while multiple imputation techniques (MICE (van2011mice), SICE (khan2020sice)) generate multiple candidate values to improve robustness. However, these approaches often rely on strong assumptions, limiting their adaptability to complex data structures.
Machine learning (ML) and deep learning (DL) have improved imputation by capturing intricate data dependencies. Neural network-based models like RNNI (rumelhart1986learning) and MLPI (garcia2010pattern) enhance generalization, while autoencoders (MIDAE (gondara2018mida), SDAE (costa2018missing)), VAEs (HI-VAE (nazabal2020handling)), and graph neural networks (GRAPE (you2020handling), IGRM (zhong2023data)) provide structural awareness. GAN-based approaches (GAIN (yoon2018gain), WGAN (arjovsky2017wasserstein)) further refine imputation by generating plausible missing values. More recently, large language models (LLMs) like GPT-3 (narayan2022can) and Table-GPT (li2024table) leverage zero-shot and few-shot learning for imputation, offering strong generalization but facing limitations in handling numerical data. Despite these advancements, DL-based methods remain computationally expensive and less interpretable than traditional approaches.
@@ 7.2. Error Detection
Error detection is essential for identifying anomalies in relational data. Traditional methods include statistical techniques (e.g., PCA (shyu2003novel), dBoost (pit2016outlier)) and constraint-based approaches enforcing integrity constraints (e.g., FDs (mandros2017discovering), CFDs (fan2008conditional), DCs (chu2013discovering)). Additionally, clustering-based methods (e.g., DBSCAN (ester1996density), OPTICS (ankerst1999optics)) and outlier detection techniques (e.g., LOF (breunig2000lof), iForest (liu2008isolation)) have also been widely used to capture anomalies.
Beyond traditional approaches, machine learning (ML)-based approaches enhance detection by integrating multiple detection signals and classifier models (e.g., Raha (mahdavi2019raha), HoloDetect (heidari2019holodetect)). Unsupervised (e.g., Uni-Detect (wang2019uni)) and active learning techniques (e.g., ED2 (neutatz2019ed2)) further improve adaptability by reducing reliance on labeled data. Deep learning (DL)-based methods advance error detection by employing autoencoders for anomaly detection (e.g., RandNet (chen2017outlier), REPEN (pang2019deep)) and utilizing semi-supervised learning to model normal and erroneous data distributions (e.g., DeepSAD (ruff2019deep)). Recently, large language models (LLMs) have been fine-tuned for tabular data anomaly detection (e.g., ADTDL (li2024anomaly)), showing potential despite computational challenges.
@@ 7.3. Error Repairing
Error repairing aims to restore data integrity by correcting erroneous values. Traditional methods include statistical reasoning (e.g., ERACER (mayfield2010eracer), SCARE (yakout2013don)) and constraint-based approaches enforcing FDs and CFDs (e.g., NADEEF (ebaid2013nadeef), Holistic (chu2013holistic)). Clustering-based strategies leverage data similarity for correction (e.g., DORC (song2015turn), DISC (song2021saving)).
Recent ML-based methods integrate probabilistic reasoning and external data sources (e.g., HoloClean (rekatsinas2017holoclean), BoostClean (krishnan2017boostclean)) to refine repairs. DL-based approaches enhance repair by learning data dependencies (e.g., DeepClean (zhang2018deepclean)) and leveraging large-scale contextual embeddings (e.g., LLMClean (biester2024llmclean)). While these methods improve repair accuracy, they remain computationally expensive, requiring further optimization (tong2016online).